If 75 litres cost £1000 what does 1 litre cost

regards
Sarah

Answer by Theo(13342)

i believe the symbol is for british pounds.
i'll assume that and you can make adjutments to whatever unit you are referring to accordingly.
75 liters cost 1000 british pounds.
1 liter will cost 1000/75 = 40/3 british pounds.
algebraically you would solve as follows:
let L represent 1 liter.
let P represent 1 pound.
your equation is:
75 * L = 1000 * P
divide both sides of this equation to gt:
L = (1000/75) * P
simplify to get:
L = (40/3) * P
that's equivalent to 13 and 1/3 pounds which has a decimal equivalent of 13.333333 pounds which you can round to the nearest decimal point of your choice.